# Pytorch-Wildlife: A Collaborative Deep Learning Framework for Conservation

### v1

@@ 1 Introduction
Deep learning technologies, such as Convolutional Neural Networks (CNNs) [15, 11], have enhanced the ability to process large quantities of data with complex structures like images. Moreover, they show remarkable performance in animal detection and classification [4, 19, 10]. Yet, implementing these technologies in practical conservation efforts presents several challenges. To effectively integrate these approaches, we identify three key aspects that require consideration: accessibility, scalability, and transparency. First, accessibility refers to the ease with which these models can be utilized; they should have easy installation processes and be compatible across different operating systems. Also, the model should provide a user interface to facilitate its use by non-technical users. Secondly, scalability focuses on the framework’s ability to add new features and to adapt to the user’s needs, ensuring its applicability across various scenarios. Finally, transparency, on the other hand, advocates for a fully open-source framework, which allows users to understand the undergoing process at each step and enables the development of new projects based on this framework. Although there are frameworks that implement some of these characteristics [6, 12, 5, 1, 26], these solutions either rely on proprietary software, do not share the code base, or have a code base that is less friendly to users with a limited engineering background and complicated for new features expansion.
@@
In this study, we introduce Pytorch-wildlife, an open-source deep-learning framework designed specifically for conservation that emphasizes accessibility, scalability, and transparency. Pytorch-wildlife can be installed via pip on any operating system that supports Python, featuring a modular architecture that facilitates the rapid incorporation of new features, models, and datasets. This framework includes a comprehensive model zoo that provides various models for animal detection and classification alongside a user interface designed for non-technical users to interact with all of its features. Pytorch-Wildlife aims to provide a unified and versatile AI platform for the conservation community and to expand to a broader range of conservation tasks, including detection, classification, and segmentation, from various data sources such as camera traps, bioacoustics, and overhead imagery. Pytorch-wildlife is completely open source, and the code is available at https://github.com/microsoft/CameraTraps.
@@ 2 Pytorch-Wildlife
3 Application in real-life scenarios
@@
In our inference of the Amazon Rainforest dataset, we implement a 98% confidence threshold as part of a human-in-the-loop procedure. We observe that the model predicts 90% of the data with recognition confidence exceeding this threshold. Furthermore, within this high-confidence subset, the model achieves an average classification accuracy of 92%. This means that, after filtering out empty images with MegaDetector, only 10% of the detected animal objects require human validation.
@@ 5 Ethical considerations
Sharing camera trap data with spatial metadata poses risks like exposing endangered species to poachers. To mitigate this, Pytorch-wildlife platform will generalize location information through a metadata filtering process. Additionally, to address privacy concerns from inadvertent human captures, human images will be removed from datasets shared through the platform.

### v2

@@ 1 Introduction
Deep learning technologies, such as Convolutional Neural Networks (CNNs) [16, 12], have enhanced the ability to process large quantities of data with complex structures like images. Moreover, they show remarkable performance in animal detection and classification [4, 21, 10]. Yet, implementing these technologies in practical conservation efforts is challenging, especially to practitioners with limited engineering backgrounds. To effectively implement these approaches, we identify three key aspects that require consideration: accessibility, scalability, and transparency. First, accessibility refers to the ease with which these models can be utilized; they should have easy installation processes and be compatible across different operating systems. Also, the model should provide a user interface to facilitate its use by non-technical users. Secondly, scalability focuses on the framework’s ability to add new features and to adapt to the user’s needs, ensuring its applicability across various scenarios. Finally, transparency, on the other hand, advocates for a fully open-source framework, which allows users to understand the undergoing process at each step and enables the development of new projects based on this framework.
In recent years, various frameworks have been developed to address the challenges associated with wildlife imagery. For instance, Trapper [5], Timelapse [11] and Camelot [13] help users to easily store, annotate, manage, map and visualize camera trap imagery. Moreover, Zooniverse [29] promotes people-powered research to help scientists to annotate large volumes of data through the creation of projects. Furthermore, platforms such as Agouti [6], Wildlife Insights [1] and DeepLab Cut [18] provide a complete set of tools that include automated wildlife recognitions through deep learning models.
However, a significant challenge remains with the interoperability of these platforms with other open-source efforts and new deep learning models. The lack of access to the source code in frameworks like Wildlife Insights and Agouti, limits the ability to expand features and develop custom tools. This is particularly relevant in a diverse field like camera trap analysis, where there is a strong demand for flexible, customizable solutions. On the other hand some open-source tools like DeepLabCut have a very specialized and narrow focus (e.g., key-point detection), making it difficult to expand their area of focus and include a wide range of tools and models.
@@
In this study, we introduce Pytorch-wildlife, an open-source deep-learning framework designed specifically for conservation that emphasizes on accessibility, scalability, and transparency. Pytorch-wildlife can be installed via pip on any operating system that supports Python, featuring a modular architecture that facilitates the rapid incorporation of new features, models, and datasets. This framework includes a comprehensive model zoo that provides various models for animal detection and classification alongside a user interface designed for non-technical users to interact with all of its features. Pytorch-Wildlife aims to provide a unified and versatile AI platform for the conservation community and to expand to a broader range of conservation tasks, including detection, classification, and segmentation, from various data sources such as camera traps, bioacoustics, and overhead imagery. Pytorch-wildlife is completely open source, and the code is available at https://github.com/microsoft/CameraTraps.
@@ 2 Pytorch-Wildlife
Model Zoo
As part of Pytorch-Wildlife’s transparency initiative, we encourage contributors to send their own deep learning models to be readily available in our model zoo. Given the unique challenges of camera trap imagery, these detection and classification models have performance limitations tied to the training data. Variables such as the species in the dataset, the geographical location, the complexity of the background, the type of climate and the resolution of the camera, among others, can significantly affect the model’s ability to predict a class.
To define the limitations of a model, we are establishing a leaderboard with standardized hidden test sets for animal detection and classification based on data from the LILA [19] dataset, this leaderboard shows the performance of the model on each hidden test, as well as a conclusion on the type of data that aligns best with the training data. Moreover, each model uploaded to the zoo through the community channel contains a ’community feedback’ component, where verified users can rate and add comments on the limitations on the model. This leaderboard will help users to select the most appropriate model for their specific requirements, ensuring better applicability and reliability in their research.
MegaDetectorV6
MegaDetectorV5 [4] is one of the golden standards in animal detection for camera trap imagery. However, its deep learning backbone is based on an older architecture, and newer detection models have achieved state-of-the-art performance on the detection task. Additionally, one of the main challenges in MegaDetectorV5 is that it is trained on the largest version of YOLOv5 [26], which makes it incompatible with smaller devices and edge computing.
To address these limitations, we trained a new deep learning model using the YOLOv9-compact [28] architecture on the MegaDetectorV5 training dataset. We call this model MegaDetectorV6-compact (MDv6-c). MDv6-c has only one-sixth of the parameters of MegaDetectorV5, making it more suitable for smaller devices.
Table 1 shows the quantitative comparison between MegaDetectorV5 (MDv5) and MDv6-c. We calculate the precision and recall of these models using the MegaDetectorV5 validation dataset to ensure a fair comparison. MDv6-c achieves a recall rate of 0.85, which is 12 percentage points higher than MegaDetectorV5. Remarkably, this improvement is accomplished with a model that has only one-sixth of the parameters of MegaDetectorV5. The higher recall on this compact-sized model allows more animals to be detected and reducing false negatives.
The compact nature of MegaDetectorV6-compact is particularly advantageous in resource-constrained environments and devices, where we can generate predictions faster and more efficient without compromising performance. In addition to improved recall, this model’s efficiency paves the way for new features in Pytorch-Wildlife that are tailored for smaller devices. These enhancements will further extend the utility and versatility of our framework, supporting practitioners in diverse field conditions.
Model Parameters Precision Recall mAP
MDv5 121M 0.96 0.73 0.85
MDv6-c 22M 0.92 0.85 0.84
Table 1: Precision and Recall comparison between MegaDetectorV5 (MDv5) and MegaDetectorV6 trained with YOLOv9-compact (MDv6-c). MDv6-c increases 12 percentage points in the recall with one-sixth of the parameters. This makes it suitable for inference in smaller devices.
In the future, we are also providing a larger version of MegaDetectorV6 using YOLOv9-extra for optimal detection performance and also a transformer based MegaDetector using RT-DETR [30] to prepare MegaDetector for the future of transformer based models.
MegaDetectorV6-compact’s pretrained weights are available in Pytorch-Wildlife.
3 Applications in real-life scenarios
@@
In our inference of the Amazon Rainforest dataset, we implement a 98% confidence threshold as part of a human-in-the-loop procedure. We observe that the model predicts 90% of the data with recognition confidence exceeding this threshold. Furthermore, within this high-confidence subset, the model achieves an average classification accuracy of 92%. This means that, after filtering out empty images with MegaDetectorv5, only 10% of the detected animal objects require human validation.
@@ 5 Ethical considerations
Sharing camera trap data with spatial metadata poses risks like exposing endangered species to poachers. To mitigate this, Pytorch-wildlife platform will generalize location information through a metadata filtering process. Additionally, to address privacy concerns, human images will be removed from datasets shared through the platform.